OpenStudy (anonymous):

Solve for n: 6(2n + 3) – 4n + 32 = 90

OpenStudy (anonymous):

12n -4n = 8n 18+32-90 = -50 8n - 50 = 0 8n = 50 n = 50/8

OpenStudy (anonymous):

Oh misscalculation its supposed to be 40*

OpenStudy (anonymous):

so n = 5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome :D